# A decade of the fast-varying ionospheric and magnetospheric magnetic fields

### v1

@@ Acknowledgements.
The mathematical formulation of this proof follows the same line as that in Appendix A. The dimension of the nullspace of the original linear system can still be identified with the that of the matrix $\sum_{k}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{S}_{k}$ (Eq. 29). The underlying assumption here is that the field coefficients can still be determined uniquely in a least squares sense in the absence of observatory bias, i.e. $\mathbf{G}_{k}^{\top}\mathbf{G}$ invertible. A corollary is that $\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{int}}$ and $\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{ext}}$ must both be full column rank, and their column spaces should be mutually linearly independent. Therefore, we can introduce a full column rank matrix $\mathbf{G}^{\prime}=\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{int}}\bm{\Lambda}-\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{ext}}$. The action of the sampling matrix $\mathbf{S}_{k}$ on $\mathbf{G}^{\prime}$ lies within the image of $\mathbf{G}_{k}$, as
$\mathbf{S}_{k}\mathbf{G}^{\prime}=\begin{pmatrix}\mathbf{S}_{k}^{\prime}\\ \mathbf{0}\end{pmatrix}(\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{int}}\bm{\Lambda}-\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{ext}})=\begin{pmatrix}\mathbf{S}_{k}^{\prime}\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{int}}\bm{\Lambda}-\mathbf{S}^{\prime}_{k}\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{ext}}\\ \mathbf{0}\end{pmatrix}=\mathbf{G}_{k}\begin{pmatrix}\bm{\Lambda}\\ -\mathbf{I}\\ \mathbf{0}\end{pmatrix}$ (36)
@@
where we used eq. (35) in the last step. As a result, the image of $\mathbf{G}^{\prime}$ lies in the nullspace of $\sum_{k}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{S}_{k}$,
@@
$\left(\sum_{k=1}^{K}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{S}_{k}\right)\mathbf{G}^{\prime}=\sum_{k=1}^{K}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{G}_{k}\begin{pmatrix}\bm{\Lambda}\\ -\mathbf{I}\\ \mathbf{0}\end{pmatrix}=\mathbf{0}.$ (37)
@@
$\mathrm{dim}\,\mathrm{ker}\left(\widetilde{\mathbf{G}}^{\top}\widetilde{\mathbf{G}}\right)=\mathrm{dim}\,\mathrm{ker}\left(\sum_{k=1}^{K}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{S}_{k}\right)\geq\mathrm{rank}(\mathbf{G}^{\prime})=\frac{1}{3}M_{0}$ (38)
@@ Appendix C Solving block-diagonal linear system with intercept
$\begin{pmatrix}\mathbf{G}_{1}&&\mathbf{0}\\ &\ddots&\\ \mathbf{0}&&\mathbf{G}_{K}\end{pmatrix}\begin{pmatrix}\mathbf{x}_{1}\\ \vdots\\ \mathbf{x}_{K}\end{pmatrix}\simeq\begin{pmatrix}\mathbf{d}_{1}\\ \vdots\\ \mathbf{d}_{K}\end{pmatrix}$ (39)
which we denote using the shorthand notation $\mathbf{G}\mathbf{x}\simeq\mathbf{d}$. Since the matrix $\mathbf{G}$ is block diagonal, the field coefficients for different time bins naturally decouple, and the least squares problem can be solved for each time bin separately. The computational and memory cost is therefore bounded by the summation of the costs for solving sub-systems $\mathbf{G}_{k}\mathbf{x}_{k}\simeq\mathbf{d}_{k}$, without involving the action of the full matrix $\mathbf{G}$.
@@
$\begin{pmatrix}\mathbf{G}&\mathbf{S}\end{pmatrix}\begin{pmatrix}\mathbf{x}\\ \mathbf{b}\end{pmatrix}=\begin{pmatrix}\mathbf{G}_{1}&&\mathbf{0}&\mathbf{S}_{1}\\ &\ddots&&\vdots\\ \mathbf{0}&&\mathbf{G}_{K}&\mathbf{S}_{K}\\ \end{pmatrix}\begin{pmatrix}\mathbf{x}_{1}\\ \vdots\\ \mathbf{x}_{K}\\ \mathbf{b}\end{pmatrix}\simeq\begin{pmatrix}\mathbf{d}_{1}\\ \vdots\\ \mathbf{d}_{K}\end{pmatrix}.$ (40)
@@
We present here an algorithm that reduces the full system into solving a series of smaller sub-systems by virtue of matrix partitioning. The algorithm consists of two steps. In the first step, the bias $\mathbf{b}$ is estimated. After removing the bias contribution from the data, the $\mathbf{x}_{k}$ vectors can then be estimated using the block-diagonal linear least squares solver in the second step. To begin with, solving the aforementioned linear least squares is equivalent mathematically to solving the normal equation
@@
$\mathbf{b}=\left(\sum_{k=1}^{K}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{S}_{k}\right)^{-1}\sum_{k=1}^{K}\mathbf{S}_{k}^{\top}\mathbf{P}_{\mathbf{G}_{k}}^{\perp}\mathbf{d}_{k}.$ (45)

### v2

@@ Acknowledgements.
The mathematical formulation of this proof follows the same line as that in Appendix A. The dimension of the nullspace of the original linear system can still be identified with the that of the matrix $\sum_{i}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{S}_{i}$ (Eq. 31). The underlying assumption here is that the field coefficients can still be determined uniquely in a least squares sense in the absence of observatory bias, i.e. $\mathbf{G}_{i}^{\top}\mathbf{G}_{i}$ invertible. A corollary is that $\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{int}}$ and $\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{ext}}$ must both be full column rank, and their column spaces should be mutually linearly independent. Therefore, we can introduce a full column rank matrix $\mathbf{G}^{\prime}=\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{int}}\bm{\Lambda}-\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{ext}}$. The action of the sampling matrix $\mathbf{S}_{i}$ on $\mathbf{G}^{\prime}$ lies within the image of $\mathbf{G}_{i}$, as
$\mathbf{S}_{i}\mathbf{G}^{\prime}=\begin{pmatrix}\mathbf{S}_{i}^{\prime}\\ \mathbf{0}\end{pmatrix}(\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{int}}\bm{\Lambda}-\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{ext}})=\begin{pmatrix}\mathbf{S}_{i}^{\prime}\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{int}}\bm{\Lambda}-\mathbf{S}^{\prime}_{i}\mathbf{G}_{0,\mathrm{obs}}^{\mathrm{ext}}\\ \mathbf{0}\end{pmatrix}=\mathbf{G}_{i}\begin{pmatrix}\bm{\Lambda}\\ -\mathbf{I}\\ \mathbf{0}\end{pmatrix}$ (38)
@@
where we used Eq. (37) in the last step. As a result, the image of $\mathbf{G}^{\prime}$ lies in the nullspace of $\sum_{i}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{S}_{i}$,
@@
$\left(\sum_{i=1}^{N_{t}}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{S}_{i}\right)\mathbf{G}^{\prime}=\sum_{i=1}^{N_{t}}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{G}_{i}\begin{pmatrix}\bm{\Lambda}\\ -\mathbf{I}\\ \mathbf{0}\end{pmatrix}=\mathbf{0}.$ (39)
@@
$\mathrm{dim}\,\mathrm{ker}\left(\widetilde{\mathbf{G}}^{\top}\widetilde{\mathbf{G}}\right)=\mathrm{dim}\,\mathrm{ker}\left(\sum_{i=1}^{N_{t}}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{S}_{i}\right)\geq\mathrm{rank}(\mathbf{G}^{\prime})=\frac{1}{3}M_{0}$ (40)
@@ Appendix C Solving block-diagonal linear system with intercept
$\begin{pmatrix}\mathbf{G}_{1}&&\mathbf{0}\\ &\ddots&\\ \mathbf{0}&&\mathbf{G}_{N_{t}}\end{pmatrix}\begin{pmatrix}\mathbf{x}_{1}\\ \vdots\\ \mathbf{x}_{N_{t}}\end{pmatrix}\simeq\begin{pmatrix}\mathbf{d}_{1}\\ \vdots\\ \mathbf{d}_{N_{t}}\end{pmatrix}$ (41)
which we denote using the shorthand notation $\mathbf{G}\mathbf{x}\simeq\mathbf{d}$. Since the matrix $\mathbf{G}$ is block diagonal, the field coefficients for different time bins naturally decouple, and the least squares problem can be solved for each time bin separately. The computational and memory cost is therefore bounded by the summation of the costs for solving sub-systems $\mathbf{G}_{i}\mathbf{x}_{i}\simeq\mathbf{d}_{i}$, without involving the action of the full matrix $\mathbf{G}$.
@@
$\begin{pmatrix}\mathbf{G}&\mathbf{S}\end{pmatrix}\begin{pmatrix}\mathbf{x}\\ \mathbf{b}\end{pmatrix}=\begin{pmatrix}\mathbf{G}_{1}&&\mathbf{0}&\mathbf{S}_{1}\\ &\ddots&&\vdots\\ \mathbf{0}&&\mathbf{G}_{N_{t}}&\mathbf{S}_{N_{t}}\\ \end{pmatrix}\begin{pmatrix}\mathbf{x}_{1}\\ \vdots\\ \mathbf{x}_{N_{t}}\\ \mathbf{b}\end{pmatrix}\simeq\begin{pmatrix}\mathbf{d}_{1}\\ \vdots\\ \mathbf{d}_{N_{t}}\end{pmatrix}.$ (42)
@@
We present here an algorithm that reduces the full system into solving a series of smaller sub-systems by virtue of matrix partitioning. The algorithm consists of two steps. In the first step, the bias $\mathbf{b}$ is estimated. After removing the bias contribution from the data, the $\mathbf{x}_{i}$ vectors can then be estimated using the block-diagonal linear least squares solver in the second step. To begin with, solving the aforementioned linear least squares is equivalent mathematically to solving the normal equation
@@
$\mathbf{b}=\left(\sum_{i=1}^{N_{t}}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{S}_{i}\right)^{-1}\sum_{i=1}^{N_{t}}\mathbf{S}_{i}^{\top}\mathbf{P}_{\mathbf{G}_{i}}^{\perp}\mathbf{d}_{i}.$ (47)